1) A 10 ppt solution is equivalent to which of the following

a)1 gram / liter

b) 0.1 grams / 100 ml

c) 10 grams / liter

d) 10 grams / 100 ml
Answer #1

ANSWER : B and C

We have, ppt = (mass of solute / mass of solution ) 1000

A) 1 gm / L

Assume density of solution = 1 g / ml , then mass of 1 L solution = 1000 g

Therefore, (1 g / 1000 g ) 1000 = 1 ppt

B) 0.1 g / 100 ml.

Assume density of solution = 1 g / ml , then mass of 100 solution = 100 g

Therefore, (1 g / 100 g ) 1000 = 10 ppt

C) 10 g per litre

Assume density of solution = 1 g / ml , then mass of 1 L solution = 1000 g

Therefore, (10 g /1000 g ) 1000 = 10 ppt

D) 10 g per 100 ml

Assume density of solution = 1 g / ml , then mass of 100 solution = 100 g

Therefore, (10 g / 100 g ) 1000 = 100 ppt